# KorNAT: LLM Alignment Benchmark for Korean Social Values and Common Knowledge

### v3

@@ Abstract
For Large Language Models (LLMs) to be effectively deployed in a specific country, they must possess an understanding of the nation’s culture and basic knowledge. To this end, we introduce National Alignment, which measures an alignment between an LLM and a targeted country from two aspects: social value alignment and common knowledge alignment. Social value alignment evaluates how well the model understands nation-specific social values, while common knowledge alignment examines how well the model captures basic knowledge related to the nation. We constructed KorNAT, the first benchmark that measures national alignment with South Korea. For the social value dataset, we obtained ground truth labels from a large-scale survey involving 6,174 unique Korean participants. For the common knowledge dataset, we constructed samples based on Korean textbooks and GED reference materials. KorNAT contains 4K and 6K multiple-choice questions for social value and common knowledge, respectively. Our dataset creation process is meticulously designed and based on statistical sampling theory and was refined through multiple rounds of human review. The experiment results of seven LLMs reveal that only a few models met our reference score, indicating a potential for further enhancement. KorNAT has received government approval after passing an assessment conducted by a government-affiliated organization dedicated to evaluating dataset quality. Samples and detailed evaluation protocols of our dataset can be found in https://selectstar.ai/ko/papers-national-alignment.
@@
[…]
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI et al., 2023) have attracted global attention due to their impressive performance and their ease of access for worldwide users. Recent research has concentrated on aligning LLMs with human values (Gabriel, 2020; Kenton et al., 2021; Ouyang et al., 2022), with the goal of ensuring LLMs behave in ways aligned with human expectations. It is, however, essential to recognize that human values and their importance are different across cultures, countries, and time periods (Davani et al., 2023; Sorensen et al., 2023). Answers that are acceptable in one culture may be entirely inappropriate in another. This becomes more important when considering that many current LLMs exhibit a bias towards English-speaking cultures (Wang et al., 2023; Zhang et al., 2023; Cao et al., 2023). Furthermore, cultural alignments have not been extensively studied in diverse cultures, as most datasets (Forbes et al., 2020; Solaiman and Dennison, 2021; Askell et al., 2021) are constructed from Western perspectives.
@@ 2 Related Works
Existing several datasets (Hendrycks et al., 2020; Forbes et al., 2020; Solaiman and Dennison, 2021) assess LMs’ basic ethics or their alignment to global values (e.g., opposition to human inequalities). However, these datasets fall short in measuring national alignment, as they solely focus on universal moral principles rather than values specific to each nation. Others (Parrish et al., 2022; Li et al., 2020; Selvam et al., 2022; Gupta et al., 2023) test social biases or stereotypes but are predominately constructed from Western perspectives. While there are efforts to reflect nation-specific social biases (Lee et al., 2023b; Jin et al., 2023; Huang and Xiong, 2023), their research is limited to social biases or stereotypes, which are insufficient to assess comprehensive national alignment. While Wang et al. (2023) investigated whether LMs incorporate opinions from multiple countries, this research is also limited as it confines its scope to only two aspects, traditional and survivals, and each question has relatively small participant responses, ranging from 10 to 20, failing to adequately represent the general opinions in the respective countries. SQuARe (Lee et al., 2023a) tests if models can keep non-toxic discussions on sensitive topics, however, it also includes few responses.
In contrast, our social value dataset differentiates itself by focusing on broader nation-specific topics not limited to biases and stereotypes, and gathering a substantial number of participant responses, an average of 219 per question. Therefore, we provide more accurate reflections of the general population’s views.
@@ 4 National Alignment Score
Assigning a single ground truth label based on the majority vote may ignore valuable information in the responses on other options (Aroyo and Welty, 2013; Cheplygina and Pluim, 2018; Davani et al., 2022). Therefore, we use the distribution of responses from the survey to measure the social value alignment. Let $r_{ij}$ be the ratio of participants choosing the $j$-th option for the $i$-th question, $q_{i}$. If a model predicts the $k$-th option for $q_{i}$, it receives an alignment score of $r_{ik}$. The final social alignment score is the average across all questions.
@@ 5.1 Experiment Settings
In our experiments, a model is prompted with an instruction (e.g., "Choose an answer from the following choices."), a question, and corresponding choices and then asked to generate a response in a zero-shot manner. For generated responses that do not exactly match with any of the choices, we employed gpt-4-1106-preview to assign the generated response to one of the choices. Considering the instability of prompting strategies (Liu et al., 2021; Min et al., 2022), we conducted experiments using five distinct yet semantically similar prompts. We tested seven models which are Llama-2 (70B) (Touvron et al., 2023), GPT-3.5-Turbo (Ouyang et al., 2022), GPT-4 (OpenAI et al., 2023), Claude-1, HyperCLOVA X [^5] https://clova-x.naver.com/ from NAVER, PaLM-2 (Anil et al., 2023), and Gemini Pro (Team et al., 2023). HyperCLOVA X is a Korean LLM extensively trained on a large Korean corpus. Prompts, post-processing of generated responses, and other additional details of experiment settings are in Appendix D.
@@ 6 Conclusion
All survey participants voluntarily reported their personal information and consented to its collection for this study. We also informed participants that their responses would be anonymized and securely protected. Additionally, we provided the option to select ‘Prefer not to answer’ for any sensitive personal information, including sexual orientation and disability, in order to respect participants’ comfort and privacy. Participants was able to discontinue the survey at any time. Compensation was adequately provided, all exceeding KRW 10,000 per hour, which surpasses the 2023 minimum wage in Korea of KRW 9,620 per hour. This study has been approved by IRB (KH2024-020).
@@
We expect that our work will considerably contribute to improving national alignment between LMs and the targeted countries, from both social values and common knowledge standpoints. Our research will inspire further studies of national alignment, fostering a more inclusive understanding and appreciation of diverse national characteristics.
Owing to the flexibility of our dataset curation framework, we strongly encourage researchers to create their own national alignment datasets. However, it is important to note that our dataset reflects the unique and regional characteristics specific to Korea. Therefore, we recommend researchers to develop their own datasets to reflect their cultural and societal contexts, making necessary adjustments to the curation process.
Since our dataset does not contain any sensitive or harmful content, we do not expect any negative ethical impacts from our research.
@@ A.2 Composition
    There are a total of 10K samples, 4K from social values and 6K from common knowledge.
@@ A.6 Distribution
    No, our dataset is intended for evaluating LLMs distributed within the country as a benchmark dataset. Only the sample dataset from the paper will be publicly available from the dataset.
@@
    Our dataset will not be distributed in any form. It is designated for use in a leaderboard system, which is operated collaboratively with the National Information Society Agency (NIA), an organization that played a pivotal role in the dataset’s construction.
@@
    Our dataset will not be distributed in any form.
@@ B.7 Participant Statistics
Table 4 shows the statistics of the 6,174 survey participants by gender, age, job, domestic area, sexual orientation, education level, annual income, religion and disability.
@@ E.1 Experiment Setting
With small LMs, we conduct experiments using likelihood-based approach. In this approach, a model is given a question accompanied by one of the candidate answers. We then compute the likelihood for each candidate, selecting the highest scoring one. We chose this method because small models tend to generate irrelevant responses. We also used five distinct yet semantically similar prompts as shown in Table 13 and Table 14. These five prompts have been carefully modified to maintain as much semantic similarity as possible. We tested four multilingual LMs and five Korean fine-tuned LMs. Korean fine-tuned models are developed upon open-sourced multilingual LMs, with additional fine-tuning on Korean corpora. For multilingual LMs, we used Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), Polyglot (Ko et al., 2023), Llama-2 (13B) (Touvron et al., 2023). For Korean fine-tuned LMs, we used KoAlpaca888beomi/KoAlpaca-llama-1-7b, KoVicuna999junelee/ko_vicuna_7b, KoAlpaca-Polyglot101010beomi/KoAlpaca-Polyglot-12.8B, KULLM-Polyglot (Lee et al., 2023c), KoLlama-2 (13B)111111beomi/llama-2-koen-13b.

### v5

@@ Abstract
For Large Language Models (LLMs) to be effectively deployed in a specific country, they must possess an understanding of the nation’s culture and basic knowledge. To this end, we introduce National Alignment, which measures an alignment between an LLM and a targeted country from two aspects: social value alignment and common knowledge alignment. Social value alignment evaluates how well the model understands nation-specific social values, while common knowledge alignment examines how well the model captures basic knowledge related to the nation. We constructed KorNAT, the first benchmark that measures national alignment with South Korea. For the social value dataset, we obtained ground truth labels from a large-scale survey involving 6,174 unique Korean participants. For the common knowledge dataset, we constructed samples based on Korean textbooks and GED reference materials. KorNAT contains 4K and 6K multiple-choice questions for social value and common knowledge, respectively. Our dataset creation process is meticulously designed and based on statistical sampling theory and was refined through multiple rounds of human review. The experiment results of seven LLMs reveal that only a few models met our reference score, indicating a potential for further enhancement. KorNAT has received government approval after passing an assessment conducted by a government-affiliated organization dedicated to evaluating dataset quality. Samples and detailed evaluation protocols of our dataset can be found in https://huggingface.co/datasets/jiyounglee0523/KorNAT.
@@
Jiyoung Lee1thanks:   Equal Contribution., Minwoo Kim2∗, Seungho Kim1, Junghwan Kim2 Seunghyun Won3, Hwaran Lee4, Edward Choi1 1KAIST AI  2SelectStar  3Seoul National University Bundang Hospital  4NAVER AI Lab 1{jiyounglee0523, shokim, edwardchoi}@kaist.ac.kr 2{mwkim, jh.kim}@selectstar.ai3shwon0213@gmail.com4hwaran.lee@navercorp.com
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI et al., 2023) have attracted global attention due to their impressive performance and their ease of access for worldwide users. Recent research has concentrated on aligning LLMs with human values (Gabriel, 2020; Kenton et al., 2021; Ouyang et al., 2022), with the goal of ensuring LLMs behave in ways aligned with human expectations. It is, however, essential to recognize that human values and their importance are different across cultures, countries, and time periods (Davani et al., 2023; Sorensen et al., 2023). Answers that are acceptable in one culture may be entirely inappropriate in another. This becomes more important when considering that many current LLMs exhibit a bias towards English-speaking cultures (Wang et al., 2023; Zhang et al., 2023; Cao et al., 2023; Havaldar et al., 2023). Furthermore, cultural alignments have not been extensively studied in diverse cultures, as most datasets (Forbes et al., 2020; Solaiman and Dennison, 2021; Askell et al., 2021) are constructed from Western perspectives.
@@ 2 Related Works
Existing several datasets (Hendrycks et al., 2020; Forbes et al., 2020; Solaiman and Dennison, 2021) assess LMs’ basic ethics or their alignment to global values (e.g., opposition to human inequalities). However, these datasets fall short in measuring national alignment, as they solely focus on universal moral principles rather than values specific to each nation. Others (Parrish et al., 2022; Li et al., 2020; Selvam et al., 2022; Gupta et al., 2023) test social biases or stereotypes but are predominately constructed from Western perspectives. While there are efforts to reflect nation-specific social biases (Lee et al., 2023b; Jin et al., 2023; Huang and Xiong, 2023), their research is limited to social biases or stereotypes, which are insufficient to assess comprehensive national alignment. Several works (Wang et al., 2023; Durmus et al., 2023; Santy et al., 2023) have focused on measuring the extent to which LLMs incorporate opinions from diverse countries, which diverges from our work which focuses on alignment with one specific country. Additionally, their datasets cannot be directly used to evaluate national alignment as the questions do not account for country-specific characteristics. For example, Durmus et al. (2023) utilized general questions from global surveys, and Santy et al. (2023) sub-sampled questions from Social Chemistry (Forbes et al., 2020) and Dynahate (Vidgen et al., 2020), which do not reflect country-specific characteristics. Questions from Wang et al. (2023) are also limited as they only reflect two aspects, traditional and survivals, and each question has relatively small participant responses, ranging from 10 to 20, failing to adequately represent the general opinions in the respective countries. SQuARe (Lee et al., 2023a) tests if models can keep non-toxic discussions on sensitive topics, however, it also includes few responses.
In contrast, our social value dataset differentiates itself by focusing on broader nation-specific topics not limited to biases and stereotypes, and gathering a substantial number of participant responses, an average of 219 per question. The dataset creation process of Santurkar et al. (2023) is similar to ours. However, our contribution comes from focusing on Korea, which is under-represented in the AI industry. Furthermore, our work is distinguished in three additional aspects. First, while questions and topics are chosen by the experts in Santurkar et al. (2023), our questions are made upon keywords extracted from monthly social conflict reports and last 12 months of news articles, ensuring they accurately reflect current Korean interests and public opinions. Our method of generating questions captures broader and more timely topics than the previous work. Second, all questions have undergone two rounds of human revisions to ensure high quality and elaborateness. Third, we applied statistical sampling theory in developing our dataset, aiming to enhance its representativeness to the best of our abilities. Therefore, we provide more accurate reflections of the general population’s views.
@@ 4 National Alignment Score
Assigning a single ground truth label based on the majority vote may ignore valuable information in the responses on other options (Aroyo and Welty, 2013; Cheplygina and Pluim, 2018; Davani et al., 2022). Therefore, we use the distribution of responses from the survey to measure the social value alignment. Let $r_{ij}$ be the ratio of participants choosing the $j$-th option for the $i$-th question, $q_{i}$. If a model predicts the $k$-th option for $q_{i}$, it receives an alignment score of $r_{ik}$. Thus, the model earns a score between 0 and 1 for each question. The final social alignment score is the average across all questions.
@@ 5.1 Experiment Settings
In our experiments, a model is prompted with an instruction (e.g., "Choose an answer from the following choices."), a question, and corresponding choices and then asked to generate a response in a zero-shot manner. For generated responses that do not exactly match with any of the choices, we employed gpt-4-1106-preview to assign the generated response to one of the choices. Considering the instability of prompting strategies (Liu et al., 2021; Min et al., 2022), we conducted experiments using five distinct yet semantically similar prompts. We tested seven models which are Llama-2 (70B) (Touvron et al., 2023), GPT-3.5-Turbo (Ouyang et al., 2022), GPT-4 (OpenAI et al., 2023), Claude-1, HyperCLOVA X (Yoo et al., 2024) from NAVER, PaLM-2 (Anil et al., 2023), and Gemini Pro (Team et al., 2023). HyperCLOVA X is a Korean LLM extensively trained on a large Korean corpus. Prompts, post-processing of generated responses, and other additional details of experiment settings are in Appendix D.
@@ 5.2.1 Quantitative Results
5.2.2 Cross-national Prompting
Cross-national Prompting (CP) (Durmus et al., 2023) is a prompting method that includes the question, ‘How would someone from [country X] respond to this question?’. We conducted experiments by replacing ‘[country X]’ with Korea and USA, respectively.
Model SVA A-SVA N-SVA
GPT-3.5-Turbo 0.290$\pm$0.008 0.435$\pm$0.016 0.315$\pm$0.003
   Korean CP 0.334$\pm$0.004 0.503$\pm$0.008 0.286$\pm$0.003
   USA CP 0.324$\pm$0.006 0.486$\pm$0.011 0.283$\pm$0.006
GPT-4 0.260$\pm$0.024 0.448$\pm$0.036 0.300$\pm$0.023
   Korean CP 0.332$\pm$0.011 0.528$\pm$0.012 0.332$\pm$0.009
   USA CP 0.309$\pm$0.016 0.455$\pm$0.024 0.377$\pm$0.009
Claude-1 0.286$\pm$0.027 0.407$\pm$0.037 0.321$\pm$0.039
   Korean CP 0.227$\pm$0.016 0.276$\pm$0.018 0.354$\pm$0.026
   USA CP 0.220$\pm$0.032 0.274$\pm$0.040 0.310$\pm$0.039
HyperCLOVA X 0.253$\pm$0.005 0.318$\pm$0.009 0.414$\pm$0.001
   Korean CP 0.332$\pm$0.020 0.505$\pm$0.032 0.299$\pm$0.004
   USA CP 0.319$\pm$0.007 0.492$\pm$0.012 0.290$\pm$0.008
Gemini Pro 0.303$\pm$0.006 0.513$\pm$0.003 0.312$\pm$0.009
   Korean CP 0.333$\pm$0.020 0.505$\pm$0.032 0.299$\pm$0.005
   USA CP 0.319$\pm$0.007 0.492$\pm$0.012 0.290$\pm$0.008
Table 2: Average and standard deviation of social value alignment using Cross-national Prompting on Final Adjustment. Bold indicates the better performance among Korean and USA CP.
Table 2 presents the social value alignment in the Final Adjustment. Both Korean an USA CP improved the alignment scores, except for Claude-1. When comparing Korean and USA CP, Korean CP generally performed well across all metrics in all models, except for one case. This suggests that the structure of prompts influences the social value alignment of LLMs.
@@ 6 Conclusion
The evaluation in a multiple-choice format is not the best choice to evaluate the models’ capabilities, as models can produce different outputs in free-form generation (Röttger et al., 2024). We could not cover thorough free-form generation evaluation since it is a relatively new and undergoing research field.
@@
All survey participants voluntarily reported their personal information and consented to its collection for this study. We also informed participants that their responses would be anonymized and securely protected. Additionally, we provided the option to select ‘Prefer not to answer’ for any sensitive personal information, including sexual orientation and disability, in order to respect participants’ comfort and privacy. Participants was able to discontinue the survey at any time. Compensation was adequately provided, all exceeding KRW 10,000 per hour, which surpasses the 2023 minimum wage in Korea of KRW 9,620 per hour. This study has been approved by KAIST IRB (KH2024-020).
@@
We expect that our work will considerably contribute to improving national alignment between LMs and the targeted countries, from both social values and common knowledge standpoints. Our research will inspire further studies of national alignment, fostering a more inclusive understanding and appreciation of diverse national characteristics. We believe that an LLM must incorporate both social values and common knowledge in order to be publicly used in a country for purposes on demand. Understanding common knowledge is equally as important as understanding social values in tasks that require such knowledge. For example, a model must understand compulsory educational knowledge in order to act as a chatbot to assist in administrative matters, or help in assisting students to learn in school. Note that for countries that share their languages (e.g., Spanish) with others, it is important to consider the linguistic capability and culture-specific alignment separately.
Owing to the flexibility of our dataset curation framework, we strongly encourage researchers to create their own national alignment datasets. However, it is important to note that our dataset reflects the unique and regional characteristics specific to Korea. Nationally well-aligned model with one country may not be generalizable to another, especially those with significant cultural differences. Therefore, we recommend researchers to develop their own datasets to reflect their cultural and societal contexts, making necessary adjustments to the curation process. Since our dataset does not contain any sensitive or harmful content, we do not expect any negative ethical impacts from our research.
@@ A.2 Composition
    There are a total of 10K samples: 4K from social values and 6K from common knowledge.
@@ A.6 Distribution
    Our dataset is intended for evaluating LLMs distributed within the country as a benchmark dataset. Only small samples from the dataset are publicly available now. However, we are planning to release the full dataset in December 2024 on AI hub.
@@
    After the leaderboard, which will be operated collaboratively with the National Information Society Agency (NIA), the full dataset will be released in December 2024 on AI hub.
@@
    The full dataset will be released in December 2024 on AI hub.
@@ B.7 Participant Statistics
Gender by Age Group
Age Group Male Female Total
20s 566 9.17% 509 8.25% 1,076 17.43%
30s 570 9.24% 527 8.54% 1,098 17.78%
40s 678 10.98% 658 10.66% 1,336 21.64%
50s 719 11.65% 713 11.54% 1,432 23.19%
60 and over 603 9.76% 630 10.20% 1,232 19.96%
Total 3,137 50.80% 3,037 49.20%
(a) Koran population by age group and gender adjusted by total number of survey participants.
Job
Agriculture, Forestry, and Fisheries 16 0.26%
Mining 5 0.09%
Manufacturing 1,646 26.66%
Electricity, Gas, Steam, and Air Conditioning Supply 34 0.55%
Water Supply; Sewerage, Waste Management, and Remediation Activities 47 0.76%
Construction 336 5.44%
Wholesale and Retail Trade 617 10.00%
Transportation and Storage 294 4.76%
Accommodation and Food Service 161 2.61%
Information and Communication 311 5.04%
Finance and Insurance 242 3.92%
Real Estate Activities 147 2.38%
Professional, Scientific, and Technical Services 533 8.63%
Business Facility Management, Business Support, and Rental Services 476 7.71%
Public Administration, Defense, and Social Security Administration - -%
Educational Services 258 4.18%
Health and Social Work Services 872 14.12%
Arts, Sports, and Recreation Related Services 59 0.95%
Associations and Organizations, Repair and Other Personal Services 119 1.93%
Activities of Households as Employers; Undifferentiated Goods- and Services-Producing Activities of Households for Own Use - -%
International and Foreign Institutions - -%
(b) Korean population grouped by their respective jobs adjusted by the total number of survey participants. There are a total of 22 groups, as defined by the Statistical classification of economic activities in the European Community (NACE). ‘-’ indicates NaN value where Korean population ratio is not found.
Domestic Area
Gyeonggi, Incheon 1,999 32.38%
Seoul 1,129 18.29%
Gyeongsang region 1,514 24.51%
Chungcheong region 668 10.82%
Jeolla region 599 9.70%
Gangwon region 184 2.98%
Jeju 81 1.32%
Education Level (Graduated or Attending)
Graduate school 376 6.09%
4-year college 2,093 33.91%
Junior college 1,022 16.56%
High school 2,153 34.87%
Middle school 385 6.23%
Elementary school 144 2.34%
(c) Korean population grouped by their domestic areas and current education levels adjusted by the total number of survey participants.
Annual Income ($\times$ KRW 1,000,000)
$\sim$ 30 1,803 29.2%
30 $\sim$ 50 1,266 20.5%
50 $\sim$ 70 994 16.1%
70 $\sim$ 100 1,006 16.3%
100 $\sim$ 1,099 17.8%
Religion
No Religion 3,461 56.06%
Protestantism 1,218 19.73%
Buddhism 959 15.33%
Catholicism 490 7.93%
Other 46 0.75%
(d) Korean population grouped by their annual incomes and religions adjusted by the total number of survey participants.
Table 6: Information of the Korean population adjusted by the total number of survey participants. For each group, we show the number of adjusted Korean population along with the percentage it represents.
Table 5 shows the statistics of the 6,174 survey participants by gender, age, job, domestic area, sexual orientation, education level, annual income, religion and disability. Table 6 presents that statistics of Korean population adjusted by the total number of survey participants.
@@ E.1 Experiment Setting
With small LMs, we conduct experiments using likelihood-based approach. In this approach, a model is given a question accompanied by one of the candidate answers. We then compute the likelihood for each candidate, selecting the highest scoring one. We chose this method because small models tend to generate irrelevant responses. We also used five distinct yet semantically similar prompts as shown in Table 15 and Table 16. These five prompts have been carefully modified to maintain as much semantic similarity as possible. We tested four multilingual LMs and five Korean fine-tuned LMs. Korean fine-tuned models are developed upon open-sourced multilingual LMs, with additional fine-tuning on Korean corpora. For multilingual LMs, we used Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), Polyglot (Ko et al., 2023), Llama-2 (13B) (Touvron et al., 2023). For Korean fine-tuned LMs, we used KoAlpaca777beomi/KoAlpaca-llama-1-7b, KoVicuna888junelee/ko_vicuna_7b, KoAlpaca-Polyglot999beomi/KoAlpaca-Polyglot-12.8B, KULLM-Polyglot (Lee et al., 2023c), KoLlama-2 (13B)101010beomi/llama-2-koen-13b.